OpenStudy (anonymous):

HELP!!!! A pair of shoes costs $30.99 and the state sales tax is 5%. Use the formula C = p + rp to find the total cost of the shoes, where C is the total cost, p is the price, and r is the sales tax rate.

OpenStudy (mathstudent55):

I just answered this question.

OpenStudy (anonymous):

not really

OpenStudy (jdoe0001):

I dont' see how the state tax comes in...but \(\large { c=p+rp\quad \begin{cases} p=30.99\\ r=30\%\to \frac{30}{100}\\ \end{cases} }\) plug and chug

OpenStudy (mathstudent55):